OpenStudy (anonymous):

Find the a volume and b surface area of this cube. Dimensions are in millimeters. Height =5 length=5 width=5

Directrix (directrix):

Volume of a cube is the cube of an edge and equals 5^3 = 125 cubic millimeters. A cube has six faces, all congruent squares of side 5. The area of one of the square faces is 5*5 = 25 square millimeters The surface area of the cube = 6 * (25) = 150 square millimeters.